The market supply curves and market demand curves for books are given as follows:
Supply curve: P = 0.000002Q Demand curve: P = 11 – 0.00002Q
The short-run marginal cost curve: MC = 0.1 + 0.0009Q
Question 1. At the above short-run equilibrium level, the firm is …
1. making a profit of R1 000 000
2. making a loss of R1 000 000
3. making zero economic profit.

Question 2
Assuming all firms in the market are identical, how many firms are producing books?
1. 5 firms
2. 50 firms
3. 500 firms
4. Given the information provided, it cannot be determined.
1
Expert's answer
2019-03-26T06:02:34-0400

1) Profit is maximized, when MR = MC.

MR = TR' = (P*Q)' = 11 - 0.00004Q,

11 - 0.00004Q = 0.1 + 0.0009Q,

0.00094Q = 10.9,

Q = 11,596 units.

At this quantity P = 11 - 0.00002*11596 = R10.77.

If we suppose, that MC = ATC, then the total profit is:

TP = (P - ATC)*Q = (10.77 - 0.1 - 0.0009*11596)*11596 = R2708.82.

But if P = ATC, then the firm is making zero economic profit.

2) Individual firm will produce such output, for which MR = MC.

MR = TR' = (P*Q)' = 11 - 0.0004Q,

11 - 0.00004Q = 0.1 + 0.0009Q,

0.00094Q = 10.9,

Q = 11,596 units.

Market is in equilibrium, when Qd = Qs or Pd = Ps, so:

0.000002Q = 11 - 0.00002Q,

0.000022Q = 11,

Q = 500,000 units.

So, the number of firms is 500,000/11,596 = 43.

That's why the nearest answer is 2. 50 firms.